OpenStudy (anonymous):

Please help, I don't know how to solve these at all. Find the first five terms of each sequence: a_n= 1/2a_n-1 , where a_1=20

OpenStudy (anonymous):

@abb0t will you please help me if you have time?

OpenStudy (amistre64):

the first term is given to you, what is the first term?

OpenStudy (amistre64):

after that, what is confusing you about the rule that they state?

OpenStudy (anonymous):

I don't understand really anything on this math unit I'm doing such as explicit and recursive sequences and that was all the information given to me for that specific problem.

OpenStudy (amistre64):

we want to find: a1 a2 a3 a4 a5 we are given the value of a1, and a rule that defines how each new term is calculated. a_n = 1/2 a_n-1 ------------------ what is a_2? let n=2 a_2 = 1/2 a_1 ------------------- what is a_3? let n=3 a_3 = 1/2 a_2 ------------------- etc ...

OpenStudy (anonymous):

So a is replaced by n?

OpenStudy (amistre64):

a is the name of a function, that function takes on the values of n. are you comfortable with the f(x) notation? if so, this is just an a(n) equivalent

OpenStudy (anonymous):

Ok so f(9) for example would be the same as n(9), correct?

OpenStudy (amistre64):

if we wanted to find the value of a function f(x), when x=9, then we would use the notation f(9) all that has changed in this case is the name of the function, and the domain that we can use. f changes to a x changes to n f(x) becomes a(n)

OpenStudy (anonymous):

Ok I see so it's basically the same type of problem just different variables, Thanks for the help!

OpenStudy (amistre64):

if you want to use f(x), then lets use it: f(x) = 1/2 f(x-1) given that f(1) = 20, what is f(2)? f(2) = 1/2 f(2-1) f(2) = 1/2 f(1) ^^ but we know f(1) = 20 therfore f(2) = 1/2 * 20 = 10

OpenStudy (amistre64):

f(3) = 1/2 f(3-1) f(3) = 1/2 f(2) ^^ but we know f(2) = 10 f(3) = 1/2 * 10 = 5 and it just goes on from there in the same way

OpenStudy (anonymous):

Ok so it decreases by intervals of 5.

OpenStudy (amistre64):

well, each new terms is half of the previous one

OpenStudy (amistre64):

20 to 10 is not a decrease of 5 :)

OpenStudy (anonymous):

I'm sorry I meant how it decreases from 20 to 10 that's a decrease of 10 then from 10 to 5 is a decrease of 5, both 10 and 5 are divisible by 5 (I think that's the correct term) lol

OpenStudy (amistre64):

thats a side trivia yes. but 5/2 is not divisible by 5 so thats not really a pattern worth sticking your guns to is all a1 = 20 a2 = half of 20 = 10 a3 = half of 10 = 5 a4 = half of 5 = 2.5 a5 = half of 2.5 = ____ well the last term we need to find.

OpenStudy (anonymous):

which is 2 and you keep plugging in the numbers to find the values?

OpenStudy (amistre64):

our rule is simple what ever the last term we found was, the next term is half of it. \[\underbrace{a_n}_{\large new~term} = 1/2 ~ \underbrace{a_{n-1}}_{\large prior~term}\]

OpenStudy (anonymous):

Ok I think I get it now and if it was a 2 instead of 1/2 the value would be 2 more than what it was originally right?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

Ok got it I believe I was just trying to overthink the problem, thanks for the help :)

OpenStudy (amistre64):

:) youre welcome.